# Studying Differential Mental Health Expressions in India

### v2

@@ Abstract
Psychosocial stressors and the symptomatology of mental disorders vary across cultures. However, current understandings of mental health expressions on social media are predominantly derived from studies in WEIRD (Western, Educated, Industrialized, Rich, and Democratic) contexts. In this paper, we analyze mental health posts on Reddit made by individuals in India, to identify variations in online depression language specific to the Indian context compared to users from the Rest of the World (ROW). Unlike in Western samples, we observe that mental health discussions in India additionally express sadness, use negation, are present-focused, and are related to work and achievement. Illness is uniquely correlated to India, indicating the association between depression and physical health in Indian patients. Two clinical psychologists validated the findings from social media posts and found 95% of the top-20 topics associated with mental health discussions as prevalent in Indians. Significant linguistic variations in online mental health-related language in India compared to ROW, emphasize the importance of developing precision-targeted interventions that are culturally appropriate.
Studying Differential Mental Health Expressions in India
**footnotetext: These authors served as co-first authors, alphabetically ordered by last name.
@@ 1 Introduction
Over 197 million individuals in India are diagnosed with mental health disorders Sagar et al. (2020), a disproportionate majority of whom do not receive treatment Singh (2018). The treatment gap for mental health disorders goes up to 95% in India, which is the highest across Asian countries and more severe compared to the gap of 78% in the United States (US) Murthy (2017); Naveed et al. (2020). Stigma, limited accessibility to mental healthcare, and a lack of awareness of mental health disorders, cumulatively contribute to diagnostic barriers for mental health care Meshvara (2002); Lahariya (2018); Krendl and Pescosolido (2020).
@@
Cognitive impairments such as self-rumination and negative evaluation owing to depression can be measured from language. For instance, increasing use of the first-person singular pronouns reflects self-rumination Spasojević and Alloy (2001); Holtzman et al. (2017), and phrases expressing self-deprecation indicate negative self-evaluation Kovacs and Beck (1978). Social media platforms are increasingly used for mental health-related conversations and seeking support in India Akbar et al. (2020). Language in social media posts can be harnessed to measure cultural differences in the manifestation of depression  Chancellor and De Choudhury (2020); Guntuku et al. (2017); Eichstaedt et al. (2018). Reddit, in particular, provides unique affordances, such as the choice to be anonymous, for individuals to openly discuss their experiences without any character limit and seek support in dedicated threads De Choudhury and De (2014); Boettcher (2021). Language-based AI models for screening depression can facilitate timely, non-intrusive targeted assessments, including support and treatment, especially in under-resourced contexts such as India, alleviating the challenges associated with traditional assessment methods  World Health Organization (2021).
@@
While it is long known that depression manifests differently across cultures Manson (1995), but, how is still an unanswered research question. Lacking awareness of cultural differences in the manifestation of depression leads to misdiagnosis even in clinical settings Bailey et al. (2009); Lewis-Fernandez et al. (2005). More recently, Rai et al. (2024) demonstrated differences in the language of White and Black individuals with depression, raising questions about the generalizability of widely acknowledged cognitive impairments across populations. This study is the first to focus on depression expressions in social media language posted by users in India. More importantly, we collaborate with clinical psychologists in India to validate trends and themes obtained from social media posts, which bolsters this computational study. A related previous study by De Choudhury et al. (2017) grouped India with other developing nations - for example, South Africa.Another study on understanding Indian population work by Pendse et al. (2019), examined help-seeking patterns rather than depression markers, to inform social media platform design.
@@
Understanding how the language of people with depression varies across cultures calls for interpretable features such as n-grams and Linguistic Inquiry and Word Count (LIWC) Boyd et al. (2022) relevant to other stakeholders such as psychologists Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Using a blend of cross-disciplinary language features and machine learning models, we address the overarching question, if and how the mental health expressions of Indian users on social media are different from the rest of the world by answering the following:
@@
  • Do the linguistic expressions of depression specific to India differ enough to reliably distinguish individuals from MH-India compared to those from MH-RoW?
This paper bridges two critical gaps from previous literature. First, the paper identifies mental health expressions specific to India, the world’s most populous country, by mining Reddit threads. Second, it engages with clinical psychologists practicing in India to validate the empirical findings, providing a culturally informed assessment of cross-country comparisons of mental health expressions.
@@ 2 Data
Reddit offers a platform for individuals to share their mental health journey and seek support anonymously, thereby making it a rich source to understand the symptomatology and sequelae of mental health De Choudhury and De (2014); Park et al. (2018); Gaur et al. (2018); Lokala et al. (2022). Other more widely used social media platforms such as Mx Takatak, Moj, Tiktok in India lack APIs for data collection oos . Previously, Reddit posts have been used for identifying shifts to suicidal ideation De Choudhury et al. (2016), depression symptoms Gaur et al. (2018); Liu et al. (2023), and the mental health expressions of immigrants Mittal et al. (2023b), among others.
Figure 1: The count of users for each country in the Rest of World control group (log scale). Demonstrates that the large majority of users in the ROW group are geolocated to Western countries. The "Others" Category contains countries with less than 10 users, including Belgium (9), Italy (9), Mexico (6), Malaysia (5), Romania (4), Croatia (4), UAE (2), South Africa (2), China (2), Spain (2), Greece (2), Denmark (1), Finland (1), Iceland (1), Japan (1), South Korea (1), Poland (1), Russia (1), Singapore (1), Thailand (1), Turkey (1) and Vietnam (1).
@@ 2.1 Subreddits: Mental Health vs Control
Most posts and comments in the dataset were posted between 2019 and 2020. As a preprocessing step, we removed deleted usernames and null messages. Assuming that users posting in India-specific subreddits are likely to be Indians, we identified India-focused subreddits (See Appendix A) and then grouped together the users who posted in these subreddits. After this step and filtering users who posted at least 500 words (excluding comments), we grouped users into four groups:
@@
The first group (MH-India) is our group of interest; the remaining are controls. We specifically chose posts as the scope of this study was to obtain expressions of experiences with mental health challenges rather than interactions with others’ mental health challenges.
@@ 2.2 Geolocation - India vs ROW
We used the geolocation inference approach Harrigian (2018) as a second layer of verification for user location. The geolocation model is a location estimation model that utilizes word usage, the frequency distribution of subreddit submissions, and the temporal posting habits of each user to determine their location. Specifically, we use the pre-trained GLOBAL inference model [^1] https://github.com/kharrigian/smgeo/tree/master#models to geolocate users in our dataset. We removed any users not geolocated to their group based on subreddit classification. For example, users in MH-India who are not geolocated to India and users in MH-RoW who are geolocated to India were removed. This functioned as a two-step verification to ensure that users in MH-India were from India. Ultimately, 1200 users out of the initial 4185 users were left in the MH-India group, and 930 users out of 2622 were left in the Control-India group. Most users in the ROW group were geolocated to the US (See Fig 1), affirming the dominance of West-centric data on Reddit.
@@ 2.3 Matching Control groups with users in MH-India
Age and gender are well-known confounders in behavioral health studies Schwartz et al. (2013). We first estimated the age and gender of every user in our dataset using a machine-learning approach described in Appendix B. We then matched the users from our group of interest, i.e., MH-India, with those in control groups (MH-ROW, and Control-ROW) on these two covariates. Owing to the smaller sample size, matching was not performed for the Control-India group. However, the age distribution across the four groups was fairly similar before matching, with the average age being 25 for the MH-India, Control-India, and Control-ROW groups and 24 for the MH-ROW group.
Ideally, the focus and control group samples should have indiscernible covariates. However, exact matching Rosenbaum (2020) is difficult to achieve without dropping a large set of samples. Coarsened Exact Matching (CEM) Iacus et al. (2009) is a softer version of Exact Matching, which stretches the matching criteria wide enough to avoid dropping samples that are similar but not an exact match. We implement CEM using MatchIt package Stuart et al. (2011) in R and set the distance to ‘Mahalanobis’ for one-to-one matching. The quality of matching was evaluated using Standard Mean Differences and Kolmogorov-Smirnov Statistics (See Appendix C). The mean age was 24.7 (sd= 3.41). The mean gender score was -0.97 (sd= 0.93), where a higher positive score indicates female. Table 1 shows the total number of posts and users in each of the four groups after CEM.
@@ 3.1 Language Features
    We extracted 1-3 grams from posts and created a normalized bag-of-words representation for each user. We filtered out 1-3 grams having point-wise mutual information (PMI)$\leq$5. N-grams reveal cultural idioms of distress Desai and Chaturvedi (2017) such as tension Weaver (2017) that are unique to a culture.
@@
    LIWC-22 is a dictionary comprising 102 word categories based on psycho-social states (e.g., Cognition, Social Processes, Affect, etc.) to measure mental states from the language. These word categories in LIWC are counted for each user, and the count is normalized by the total number of 1-grams for each user, thereby representing each user as a vector of 102 normalized psychosocial categories.
@@
    We used Latent Dirichlet Allocation (LDA) Blei et al. (2003) to extract latent topics in users’ timeline data. While recent neural topic modeling methods such as BERTopic Grootendorst (2022) show superior predictive accuracy, LDA has been shown to provide qualitatively robust topics Dixon et al. (2022).
    We generated three sets of topics by setting the number of topics $=[200,500,2000]$. We evaluated the topics’ quality using Topic Uniqueness (TU) Nan et al. (2019). TU represents the number of times a set of keywords is repeated across topics; a higher TU corresponds to a rarely repeated word, indicating that topics are diverse, which is favorable. Additionally, three co-authors independently reviewed the quality of topics. We set the number of topics to 2000 based on the automated (See Table A3) and manual evaluation. High granularity in topics is desired as it reveals subtle variations in language across cultures.
@@ 3.2 Correlation Analysis
To measure the association between language and the groups (i.e., MH-India vs MH-ROW), we performed ordinary least squares regressions with the feature sets. We calculated Pearson $r$ to measure the association of each feature to each group in a one-vs-all setting. This is important in supporting an understanding of how each feature set correlates with the presence of mental health challenges in India. p-values were corrected using Benjamini-Hochberg correction for multiple hypothesis testing. 102 word categories for LIWC, 2000 for LDA topics, and 23,344 1-3 grams were considered for p-value correction.
@@ 3.3 Predictive Model
To examine whether the linguistic expressions of the MH-India group sufficiently differ to discriminate it from other groups, we trained ‘one vs rest’ logistic regression models in a 10-fold cross-validation setting Rifkin and Klautau (2004). More sophisticated methods (such as XGBoost) could potentially provide higher performance, but the focus of the study is not to achieve state-of-the-art performance for group prediction but to test if sufficient discriminating evidence exists across groups to warrant culturally aligned depression models. We report the Area Under the Receiver Operating Curves (AUC) for each feature for the MH-India and MH-ROW groups.
@@ 4.1.1 N-grams
Sixty-one 1-3 grams were significantly ($p<0.05$) correlated with the MH-India group, and 156 were correlated with the MH-ROW group. Figure 2 illustrates the top 25 1-3 grams arranged in decreasing order of Pearson $r$ for both groups. First-person singular pronouns (‘i am’, ‘i’), and mentions of depression (‘depressed’, ‘depression’), are prevalent in both groups. Help-seeking phrases (‘suggest’, ‘advice’, ‘talk’, ‘please help’) are prevalent in MH-India, whereas MH-ROW discusses feelings, use clinical language (‘diagnosed’, ‘medication’) and express negative emotions (‘anxiety’, ‘bad’, ‘worse’) associated with depression. Interestingly, the MH-India group simultaneously discusses the feelings of loneliness and social relationships with parents and friends. Academic-related stress (‘college’, ‘exam’, ‘study’) is exclusively seen in discussions of the MH-India group. This is particularly interesting, considering users in both groups were matched for age, yet discussions around student-life challenges are prevalent only in MH-India.
Overall, the discussion in MH-India subreddits is centered around seeking help, whereas negative feelings are more commonly discussed in the MH-ROW group.

### v3

# Cross-Cultural Differences in Mental Health Expressions on Social Media

@@ Abstract
Culture moderates the way individuals perceive and express mental distress. Current understandings of mental health expressions on social media, however, are predominantly derived from WEIRD (Western, Educated, Industrialized, Rich, and Democratic) contexts. To address this gap, we examine mental health posts on Reddit made by individuals geolocated in India, to identify variations in social media language specific to the Indian context compared to users from the Rest of the World (RoW). Our experiments reveal significant psychosocial variations in emotions (sadness in India vs anxiety in RoW), temporal orientation (present-focused in India vs past-focused in the West), and sociocultural aspects (substance use vs work/achievement). Clinical psychologists practicing in India validated the findings and underlined significant overlap in mental health-related concerns observed in social media posts and in-person sessions. This study demonstrates the potential of social media platforms for identifying cross-cultural differences in mental health struggles (e.g. seeking help in India vs seeking peer support in RoW). Future research should investigate how mental health assessment can be culturally adapted to personalize interventions, ensuring equitable mental health care for individuals from all cultural backgrounds.
Cross-Cultural Differences in Mental Health Expressions on Social Media
**footnotetext: The authors are co-first authors.
@@ 1 Introduction
Over 197 million individuals in India are diagnosed with mental health disorders Sagar et al. (2020), a disproportionate majority of whom do not receive treatment Singh (2018). The treatment gap for mental health disorders goes up to 95% in India, which is the highest across Asian countries and more severe compared to the gap of 78% in the United States (US) Murthy (2017); Naveed et al. (2020). Stigma against mental disorders, limited accessibility to affordable mental healthcare, and a lack of awareness of mental health disorders, cumulatively contribute to diagnostic barriers for mental health care Meshvara (2002); Lahariya (2018); Krendl and Pescosolido (2020).
Figure 1: Study overview to investigate the linguistic expressions and topics specific to Indian users posting in Mental Health (MH) related subreddits (See Appendix A) compared to users across the rest of the world (RoW).
@@
A recent study reported that 64% of Indian participants used social media for mental health communication Patki and Iyer (2022). Social media language can reveal risk for mental disorders, through patterns in word choice and emotional tone Chancellor and De Choudhury (2020); Guntuku et al. (2017); Eichstaedt et al. (2018). For instance, increasing use of the first-person singular pronouns reflects excessive self-focus Spasojević and Alloy (2001); Holtzman et al. (2017), and phrases expressing self-deprecation indicate negative self-evaluation Kovacs and Beck (1978). Language-based AI models for screening mental health risks can facilitate timely, non-intrusive targeted assessments, to enable timely support and treatment, especially in under-resourced contexts such as India, alleviating the challenges associated with administering traditional assessment methods such as PHQ-9  World Health Organization (2021).
Mental disorders manifest differently across cultures Manson (1995), however, how is still an unanswered research question. Rai et al. (2024) demonstrated linguistic differences in Facebook posts of White and Black individuals with depression, raising questions about the generalizability of widely acknowledged cognitive impairments across populations. In parallel, there is growing concern about the AI models performance for predicting mental health risks in persons of color Rai et al. (2024); Aguirre et al. (2021). AI systems lacking awareness of cross-cultural differences in the manifestations of mental disorders may lead to misdiagnosis Bailey et al. (2009); Lewis-Fernandez et al. (2005).
@@
In this paper, we examine how mental health expressions specific to Indian context vary compared to the rest of the world. Previously, De Choudhury et al. (2017) examined linguistic patterns in the online language of individuals with depression but grouped India with other developing nations - for example, South Africa. Another related study on understanding Indian population work by Pendse et al. (2019), examined help-seeking patterns rather than mental health markers, to inform social media platform design. Our study is the first to focus on mental health-related expressions in social media language posted by users in India. Additionally, we collaborate with clinical psychologists in India to interpret linguistic trends and themes obtained from social media posts, which bolsters this computational study.
@@
Understanding how the language of people facing mental distress varies across cultures calls for interpretable features comprehensible to stakeholders such as psychologists and policymakers Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Psychosocial word categories (e.g., Linguistic Inquiry and Word Count (LIWC)) and topic modeling (word clusters derived using LDA) are thus preferred approaches to examine variations in language correlated with depression Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Using a blend of cross-disciplinary language features and machine learning models, we address the overarching question, if and how the mental health expressions of Indian users on social media are different from the rest of the world by answering the following:
@@
  • Do the linguistic expressions of depression specific to India differ significantly to reliably distinguish individuals from MH-India compared to those from MH-RoW?
This paper bridges two critical gaps from previous literature. First, the paper identifies mental health expressions specific to India, the world’s most populous country, by mining Reddit threads. Second, it engages with clinical psychologists practicing in India to validate the empirical findings, providing a culturally informed assessment of cross-cultural comparisons of mental health expressions.
@@ 2 Data
Reddit offers a platform for individuals to share their mental health journey and seek support anonymously, thereby making it a rich source to understand the symptomatology and sequelae of mental health De Choudhury and De (2014); Park et al. (2018); Gaur et al. (2018); Lokala et al. (2022). Previously, Reddit posts have been used for identifying shifts to suicidal ideation De Choudhury et al. (2016), depression symptoms Gaur et al. (2018); Liu et al. (2023), and the mental health expressions of immigrants Mittal et al. (2023b), among others. Other more widely used social media platforms such as Mx Takatak, Moj, Tiktok in India could be potential alternatives to collect India-specific mental health communication however, they lack APIs for data collection Oosga .
Figure 2: The count of users for each country in the Rest of World control group (log scale). Majority of users in the RoW group are geolocated to Western countries. The "Others" Category contains countries with less than 10 users, including Belgium (9), Italy (9), Mexico (6), Malaysia (5), Romania (4), Croatia (4), UAE (2), South Africa (2), China (2), Spain (2), Greece (2), Denmark (1), Finland (1), Iceland (1), Japan (1), South Korea (1), Poland (1), Russia (1), Singapore (1), Thailand (1), Turkey (1) and Vietnam (1).
@@ 2.1 Subreddits: Mental Health vs Control
Most posts and comments in the dataset were posted between 2019 and 2020. As a preprocessing step, we removed deleted usernames and null messages. Assuming that users posting in India-specific subreddits are likely to be Indians, we identified India-focused subreddits (See Appendix A) and then grouped together the users who posted in these subreddits. We considered users who posted at least 500 words (excluding comments) to ensure sufficient linguistic richness in users’ language for further analysis. We grouped users into four groups:
@@
The first group (MH-India) is our group of interest; the remaining are controls. We limit the scope of this study to Reddit posts to obtain expressions of experiences with mental health challenges rather than interactions (as in comments) with others’ mental health challenges.
@@ 2.2 Geolocation - India vs RoW
We used the geolocation inference approach Harrigian (2018) as a second layer of verification for user location. The geolocation model is a location estimation model that utilizes word usage, the frequency distribution of subreddit submissions, and the temporal posting habits of each user to determine their location. Specifically, we use the pre-trained GLOBAL inference model***https://github.com/kharrigian/smgeo/tree/master#models to geolocate users in our dataset. We removed any users not geolocated to their group based on subreddit classification. For example, users in MH-India who are not geolocated to India and users in MH-RoW who are geolocated to India were removed. This functioned as a two-step verification to ensure that users in MH-India were from India. Ultimately, 1200 users out of the initial 4185 users were left in the MH-India group, and 930 users out of 2622 were left in the Control-India group. The majority of the users in the RoW group were geolocated to the US, UK, Canada, Germany, Australia, and Sweden (See Fig 2), affirming the dominance of West-centric data on Reddit. There are hardly 100 users from non-Western nations. From now on, we use the terms RoW and Western nations interchangeably.
@@
Figure 3: Differences in Covariates before and after CEM for groups “Control-RoW” and “MH-RoW”. A control group is considered balanced with the treatment group if the difference is close to zero. Matching was not performed for Control-India due to smaller sample size.
@@ 2.3 Matching Control groups with users in MH-India
Age and gender are well-known confounders in behavioral health studies Schwartz et al. (2013). We first estimated the age and gender of every user in our dataset using a machine-learning approach (see Appendix B for method and evaluation). We then matched the users from our group of interest, i.e., MH-India, with those in control groups (MH-RoW, and Control-RoW) on these two covariates. Owing to the smaller sample size, matching was not performed for the Control-India group. However, the age distribution across the four groups was fairly similar before matching, with the average age being 25 for the MH-India, Control-India, and Control-RoW groups and 24 for the MH-RoW group.
Ideally, the focus and control group samples should have indiscernible covariates. However, exact matching Rosenbaum (2020) is difficult to achieve without dropping a large set of samples. Coarsened Exact Matching (CEM) Iacus et al. (2009) is a softer version of Exact Matching, which stretches the matching criteria wide enough to avoid dropping samples that are similar but not an exact match. We implement CEM using MatchIt package Stuart et al. (2011) in R and set the distance to ‘Mahalanobis’ for one-to-one matching. The quality of matching was evaluated using Standard Mean Differences and Kolmogorov-Smirnov Statistics (See Fig. 3). The mean age was 24.7 (sd= 3.41). The mean gender score was -0.97 (sd= 0.93), where a higher positive score indicates female. Table 1 shows the total number of posts and users in each of the four groups after matching.
@@ 3.1 Language Features
    We extracted 1-3 grams from posts and created a normalized bag-of-words representation for each user. We filtered out 1-3 grams having point-wise mutual information (PMI)$\leq$5. N-grams may reveal prevalent cultural idioms of distress Desai and Chaturvedi (2017) such as tension Weaver (2017) that are unique to a culture.
@@
    LIWC-22 is a closed dictionary comprising 102 word categories based on psycho-social states (e.g., Cognition, Social Processes, Affect, etc.) to measure mental states from the language. These word categories in LIWC are counted for each user, and the count is normalized by the total number of 1-grams for each user, thereby representing each user as a vector of 102 normalized psychosocial categories.
@@
    We used Latent Dirichlet Allocation (LDA) Blei et al. (2003) to extract latent topics in users’ timeline data to capture themes behind mental distress specific to a culture. While recent neural topic modeling methods such as BERTopic Grootendorst (2022) show superior predictive accuracy, LDA has been shown to provide qualitatively robust topics Dixon et al. (2022).
    We generated three sets of topics by setting the number of topics $=[200,500,2000]$. We evaluated the topics’ quality using Topic Uniqueness (TU) Nan et al. (2019). TU represents the number of times a set of keywords is repeated across topics; a higher TU corresponds to a rarely repeated word, indicating that topics are diverse, which is favorable. Additionally, three co-authors independently reviewed the quality of topics. We set the number of topics to 2000 based on the automated (See Table 2) and manual evaluation. Human annotators preferred high topic granularity because it highlights subtle linguistic variations across cultures.
Type #Topics TU
Posts Only 200 0.7675
Posts Only 500 0.7162
Posts Only 2000 0.7694
Table 2: Topic Uniqueness (TU) Scores for different numbers of topics.
@@ 3.2 Correlation Analysis
To measure the association between language and the groups (i.e., MH-India vs MH-RoW), we performed ordinary least squares regressions with the feature sets. We calculated Pearson $r$ to measure the association of each feature to each group in a one-vs-all setting. This is important in supporting an understanding of how each feature set correlates with the presence of mental health challenges in India. p-values were corrected using Benjamini-Hochberg correction for multiple hypothesis testing. 23,344 1-3 grams, 102 word categories for LIWC, 2000 LDA topics, were considered for p-value correction.
@@ 3.3 Predictive Model
To examine whether the linguistic expressions of the MH-India group sufficiently differ from control groups, we trained ‘one vs rest’ logistic regression models in a 10-fold cross-validation setting Rifkin and Klautau (2004). More sophisticated methods (such as XGBoost) could potentially provide higher performance, but the focus of the study is not to achieve state-of-the-art performance for group prediction but to test if sufficient discriminating evidence exists across groups to warrant culturally aligned models for estimating mental health risks. We report the Area Under the Receiver Operating Curves (AUC) for each feature for the MH-India and MH-RoW groups.
@@ 4.1.1 N-grams
Sixty-one 1-3 grams were significantly ($p<0.05$) correlated with the MH-India group, and 156 were correlated with the MH-RoW group. Figure 4 illustrates the top 30 1-3 grams arranged in decreasing order of Pearson $r$ for MH-India. Introductory phrases (i am, from india), negation (i don’t know), and help-seeking phrases (suggest, please help) are exclusively correlated with MH-India, reflecting the struggle of users in MH-India for accessing mental help. Academic-related stressors (college, exam, study) are exclusively seen in discussions in MH-India. This is particularly interesting, considering users in both groups were matched for age, yet discussions around student-life challenges are prevalent exclusively in MH-India. Self-referential pronouns such as I and me, negative feelings and symptoms mentions are more frequently seen in MH-RoW (see Fig. A1).
Figure 5: Pearson r for top-20 LIWC categories (with top-3 words) correlated with MH-India with corresponding correlation with MH-RoW. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Sadness, social behaviors, and present focus are the most prevalent categories.
Figure 6: Pearson r for top-20 LIWC categories (with top-3 words) correlated with MH-RoW with corresponding correlation with MH-India. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Substances, 1st person sing. pronouns, and Mental are the most prevalent categories.